OpenStudy (anonymous):

Simplify k-3/2 + 3k+1+3k+1/2.

OpenStudy (anonymous):

\[\frac{k-3}{2} + 3k + 1 + \frac{3k+1}{2}\] You need to make denominators equal first...

OpenStudy (anonymous):

\[\frac{k-3}{2} + \frac{3k + 1}{1} + \frac{3k+1}{2} \implies \frac{k-3}{2} + \frac{2(3k + 1)}{2} + \frac{3k+1}{2}\] \[\frac{k -3 + 6k + 2 + 3k + 1}{2} = \textsf{Solve}\]